OpenStudy (anonymous):

what is -6/4 multiplied by negative 1/4. Both are fractions

OpenStudy (jit4won):

- 6/4 X -1/4 = + 6/24 or 3/8

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

the product of two fractions is the product of the numerators over the product of the denominators. \[\frac{a}{b}*\frac{c}{d} = \frac{a*c}{b*d}\]

OpenStudy (anonymous):

then simplify of course